A small trailer has mass 400 kg. The wheels give friction, with coefficient of friction being 0.016. How much force must be exerted through the hitch to keep the trailer moving at steady speed? How much force in total must be exerted to accelerate the trailer at 1.5 m/s2?
1
Expert's answer
2019-12-02T10:06:29-0500


When moving at a constant speed, the equation of motion has the form

"F-F_{fr}=0"

where will we write

"F=F_{fr}=\mu \cdot N=\mu \cdot m \cdot g=0.016 \cdot 400 \cdot 9.81=62.784[N]"

При движении с ускорение , уравнение движения имеет вид

"F-F_{fr}=m \cdot a"

where will we write

"F=m \cdot a+F_{fr}=m \cdot a+\mu \cdot N=m \cdot a+\mu \cdot m \cdot g=400 \cdot 1.5+0.016 \cdot 400 \cdot 9.81=662.78[N]"